A student applies for two different scholarships. The probability of receiving the first scholarship is 0.3 and the probability of receiving the second is 0.4. The decisions are made independently. Find the probability the student receives exactly one scholarship.
Answer by stanbon(75887)

A student applies for two different scholarships. The probability of receiving the first scholarship is 0.3 and the probability of receiving the second is 0.4. The decisions are made independently. Find the probability the student receives exactly one scholarship.
----------------------
P(receive both) = 0.3*0.4 = 0.12
P(receive none) = 0.7*0.6 = 0.42
---------
P(receive exactly one) = 1-(0.12 + 0.42) = 0.46
========================
Cheers,
Stan H.